OpenStudy (anonymous):

The area A = πr2 of a circular water ripple changes with the radius. At what rate does the area change with respect to the radius when r = 4ft?

OpenStudy (anonymous):

I dont really know how to start this problem. Please help!

OpenStudy (anonymous):

A = pi * r^2 dA/dr = 2pi * r r = (dA/dr)/2pi = answer to a dA/dr = 2pi * 500 = 1000 * pi

OpenStudy (anonymous):

did you take the derivative of the area?

OpenStudy (anonymous):

Yes, I did

OpenStudy (anonymous):

Where di the 500 come from?